solving using the eliminations method.
8x-5y=-9
3x+5y=-2

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

add the equations
11x=-11

Answer by stanbon(75887)

solving using the eliminations method.
8x-5y=-9
3x+5y=-2
-----
Add the equations to eliminate the y-terms.
11x = -11
x = -1
----
Substitute into either equation to solve for "y":
3(-1)+5y = -2
-3 + 5y = -2
5y = 1
y = 1/5
==========
Cheers,
Stan H.